# Functional-level Uncertainty Quantification for Calibrated Fine-tuning on LLMs

### v1

@@ 6 Discussion & Conclusion
The limitation of UQ4CT lies in its dependency on predictive correctness. For general language modeling tasks such as chat completion, there lacks a clear metric on whether the response is correct or not. This limits the application of UQ4CT as naively token matching is a poor indicator of semantic correctness due to the ambiguous nature of language. For future work, we plan to continue exploring ways to adapt UQ4CT on open-ended problems that lacks a definitive metric.
Reproducibility Statement
We share our experimental details in Appendix A.1, and also provide the code and model weights for running experiments in the supplementary materials to reproduce our model performance results.
@@ Appendix A Appendix
A.1 Training Details
@@
A.2 Expected Calibration Error
@@
A.3 MMLU Distribution Shift Dataset Composition

### v2

@@ 6 Discussion & Conclusion
The limitation of UQ4CT lies in its dependency on predictive correctness. For general language modeling tasks such as chat completion, there lacks a clear metric on response correctness. This limits the application of UQ4CT as naively token matching is a poor indicator of semantic correctness due to the ambiguous nature of language. For future work, we are exploring ways to adapt UQ4CT to open-ended problems that lack a definitive optimization objective.
@@ Appendix A Appendix
A.1 Load Balancing Loss
We follow the load balancing loss in [Li et al., 2024]. Given $N$ experts indexed by $i=1$ to $N$ and a batch $B$ with $T$ tokens, the auxiliary loss is computed as:
$\mathcal{L}_{aux}=a\cdot N\cdot\sum^{N}_{i=1}\mathcal{F}_{i}\cdot\mathcal{P}_{i},$ (16)
where
$\mathcal{F}_{i}=\frac{1}{T}\sum_{x\in B}\mathbbm{1}\{argmax_{k}\mathcal{R}(x)_{k}=i\},\mathcal{P}_{i}=\frac{1}{T}\sum_{x\in B}\mathcal{R}(x)_{i}.$ (17)
Here, $\mathcal{R}(\cdot)$ is the top-k router, $\mathcal{F}_{i}$ is the fraction of tokens dispatched to expert $i$ and $\mathcal{P}_{i}$ is the fraction of the router probability allocated for expert $i$. The final loss is multiplied by the expert count $N$ to keep the loss constant as the number of experts varies, and the constant term $a$ is set to $10^{-2}$as a multiplicative coefficient, which is large enough to ensure load balancing while remaining small enough not to overwhelm the primary objective.
A.2 Experimental Results with Mistral-7B
In this section, we present the results using Mistral-7B [Jiang et al., 2023], a different decoder-based LLM backbone. Table 5 shows the results of fine-tuning Mistral-7B on $4$ common-sense reasoning tasks and one domain-specific climate question-answering task.
For each of the tasks, UQ4CT effectively calibrates the parameter mixtures, leading to the best ECE performance in $4$ out of $5$ tasks. This indicates the robustness of UQ4CT across different LLMs.
Table 5: Performance comparison of different methods fine-tuned with Mistral-7B across $4$ common sense reasoning tasks and a domain-specific task. UQ4CT shows significant ECE improvements while maintaining high accuracy.
Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
ACC $\uparrow$ LoRA $70.3_{0.62}$ $84.8_{0.47}$ $70.2_{0.84}$ $82.8_{0.62}$ $72.5_{1.6}$
MC Drop $69.6_{1.07}$ $84.6_{0.91}$ $69.6_{0.76}$ $82.6_{0.71}$ $72.5_{1.6}$
Ensemble $71.8_{1.29}$ $84.2_{0.66}$ $71.0_{1.41}$ $82.5_{0.6}$ $72.9_{2.88}$
LA $70.7_{1.82}$ $82.4_{2.05}$ $68.5_{3.31}$ $82.5_{0.77}$ $71.6_{1.56}$
MixLoRA $73.1_{0.38}$ $85.5_{1.27}$ $71.2_{1.75}$ $83.3_{1.14}$ $72.0_{1.69}$
UQ4CT $73.6_{0.28}$ $85.9_{0.82}$ $74.4_{0.82}$ $83.7_{1.22}$ $73.2_{1.29}$
ECE $\downarrow$ LoRA $10.17_{0.24}$ $9.46_{1.62}$ $18.42_{1.91}$ $13.3_{0.25}$ $13.72_{2.62}$
MC Drop $10.62_{0.51}$ $8.91_{1.35}$ $18.38_{1.66}$ $13.3_{0.31}$ $13.72_{2.61}$
Ensemble $8.72_{1.13}$ $8.72_{1.49}$ $17.0_{0.97}$ $9.14_{2.82}$ $12.86_{1.78}$
LA $5.33_{2.16}$ $20.3_{5.7}$ $21.27_{4.15}$ $\mathbf{6.41_{3.22}}$ $14.64_{2.21}$
MixLoRA $8.81_{1.03}$ $8.16_{0.99}$ $15.51_{3.86}$ $10.53_{1.73}$ $14.05_{3.09}$
UQ4CT $\mathbf{3.07_{0.83}}$ $\mathbf{5.7_{0.69}}$ $\mathbf{7.04_{0.58}}$ $\underline{7.92_{1.14}}$ $\mathbf{11.4_{1.14}}$
A.3 Incremental Weighting on Calibration Term
A.4 Deciding Number of Active Experts
One important aspect of the LoRA MoE architecture is how many experts to activate. Here, we investigate the performance impact of different number of active LoRA experts. We evaluate the model performance with $1$ to $5$ active experts with $8$ in total.
Table 6: Performance comparison of UQ4CT with varying number of experts on OBQA dataset. Top-2 expert selection strategy grants best accuracy and calibration.
Top-K ACC $\uparrow$ ECE $\downarrow$
Top-1 $74.8_{0.62}$ $7.69_{1.96}$
Top-2 $\mathbf{77.3_{1.36}}$ $\mathbf{5.0_{1.15}}$
Top-3 $75.2_{0.8}$ $5.8_{0.81}$
Top-4 $75.8_{0.53}$ $7.67_{0.46}$
Top-5 $75.3_{0.5}$ $6.29_{0.61}$
As shown in Table 6, $2$ active experts give the optimal performance in terms of accuracy and ECE scores. One expert alone cannot capture complicated functional relationships, while more than 2 experts could potentially introduce redundant functional bases to the model, which deviates the output distribution more from data distribution, thus worsening predictive and calibration performance. Additionally, more active experts lead to a more flattened distribution across experts, which hardens the alignment of parameter mixtures during fine-tuning.
A.5 Training Details
@@
A.6 Expected Calibration Error
@@
A.7 MMLU Distribution Shift Dataset Composition